OpenStudy (anonymous):

⎰((56x^6+4x^3-2x^5)dx) can be written as ax^p+bx^q+cx^r+c

OpenStudy (unklerhaukus):

\[\int 56x^6+4x^3-2x^5\,\mathrm dx\] use the sum rule \[=56\int x^6\,\mathrm dx+4\int x^3\,\mathrm dx-2\int x^5\,\mathrm dx\] then the power rule \[\boxed{\displaystyle\int x^n\,\mathrm dx=\frac1nx^{n+1}+\color{gray}c}\] \[=\]

OpenStudy (unklerhaukus):

combine all the constant of integration, in to a single constant then set your answer equal to \[ax^p+bx^q+cx^r+c\] hence, by direct comparison , solve for \(a,p,b,q,r\) and \(c\)